# Characteristics and prevalence of fake social media profiles with AI-generated faces

### v1

@@ Abstract
Recent advancements in generative artificial intelligence (AI) have raised concerns about their potential to create convincing fake social media accounts, but empirical evidence is lacking. In this paper, we present a systematic analysis of Twitter(X) accounts using human faces generated by Generative Adversarial Networks (GANs) for their profile pictures. We present a dataset of 1,353 such accounts and show that they are used to spread scams, spam, and amplify coordinated messages, among other inauthentic activities. Leveraging a feature of GAN-generated faces — consistent eye placement — and supplementing it with human annotation, we devise an effective method for identifying GAN-generated profiles in the wild. Applying this method to a random sample of active Twitter users, we estimate a lower bound for the prevalence of profiles using GAN-generated faces between 0.021% and 0.044% — around 10K daily active accounts. These findings underscore the emerging threats posed by multimodal generative AI. We release the source code of our detection method and the data we collect to facilitate further investigation. Additionally, we provide practical heuristics to assist social media users in recognizing such accounts.
@@ 1 Introduction
Recent advances in generative artificial intelligence (AI) have resulted in remarkable competencies across various domains (Cao et al., 2023). Large language models, such as ChatGPT,111chat.openai.com can understand human language and generate realistic text in different contexts (Jakesch et al., 2023; Ye et al., 2023). Generative vision models, such as Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) and diffusion models (Song and Ermon, 2019; Ho et al., 2020), produce photorealistic images that people can no longer distinguish from human-generated images (Lago et al., 2021; Nightingale and Farid, 2022; Bray et al., 2023). State-of-the-art multimodal models, such as Google’s Gemini models,222blog.google/technology/ai/google-gemini-ai can even understand visual and audio inputs and respond with text or images. This evolving landscape of generative AI models underscores a significant leap in AI technology, opening new avenues for research and applications.
These advancements also present opportunities for misuse. For instance, adversarial actors can exploit AI models to create and disseminate disinformation or conduct influence campaigns (Yamin et al., 2021; Guembe et al., 2022; Goldstein et al., 2023), threatening the integrity of online information exchange. In this study, we focus on the fake social media accounts that leverage generative AI for deceptive purposes (Bommasani et al., 2021). In documented cases, individuals have used AI-generated human images to fabricate fake personas online, leading to deception and targeted attacks. [^3] See cnn.com/2020/02/28/tech/fake-twitter-candidate-2020/ and cnn.com/2020/02/20/tech/fake-faces-deepfake/ Major social media platforms, such as Facebook444wired.com/story/facebook-removes-accounts-ai-generated-photos and LinkedIn,555npr.org/2022/03/27/1088140809/fake-linkedin-profiles have already encountered groups of fake accounts that utilize AI-generated faces.
@@
Although there are other generative vision models, we focus on GANs for several reasons. Firstly, the introduction of GAN-based models, particularly the StyleGAN series (Karras et al., 2019, 2020), marks a major milestone in the field of generative AI models (Cao et al., 2023) because they are capable of producing photorealistic images almost indistinguishable from real ones (Nightingale and Farid, 2022). Secondly, GANs are easily accessible. With their source code publicly available online, [^7] See github.com/NVlabs/stylegan and github.com/NVlabs/stylegan2 adversarial actors can easily produce fake human faces as needed. Additionally, for those without technical expertise, websites like thispersondoesnotexist.com offer an easy means to download GAN-generated images in large quantities. Last but not least, GAN-generated faces share some distinctive features, allowing us to identify them with high confidence.
In this study, we create a novel dataset of 1,353 inauthentic Twitter accounts, each utilizing a GAN-generated face as its profile picture. Upon close examination, these accounts typically feature human-like names suggesting a deliberate effort to mimic real users. These accounts engage in inauthentic activities, such as disseminating spam, executing scams, and amplifying messages in a coordinated fashion, often through automation. We further estimate the prevalence of such accounts on the platform. Leveraging the feature that GANs consistently place eyes at the same locations, together with human annotation, we develop a simple yet effective method to identify GAN-generated profiles in the wild. Applying this method to a random sample of 254,275 active Twitter users, our analysis indicates a lower bound for the prevalence of accounts using GAN-generated faces between 0.021% and 0.044%.
@@ 2.1 GAN-generated face detection
Although we also need to detect GAN-generated faces in this study, our case differs from previous studies significantly. The aforementioned methodology papers typically aim to distinguish generated and real human faces from datasets such as FFHQ (Karras et al., 2019) and CelebA (Liu et al., 2015). On the other hand, our objective is to identify generated faces among random Twitter profiles, which presents unique challenges. Firstly, only a subset of Twitter profiles feature human faces, and their styles may differ greatly from GAN-generated ones (see Fig. 1(c) for examples). Secondly, the instances of GAN-generated faces may be relatively scarce, necessitating a detection method with high accuracy to address the class imbalance issue and provide a reliable estimate of the prevalence of fake accounts. Given these differences, existing methods are not suitable for our purpose.
@@ 2.2 GAN-generated faces on social media
Unfortunately, these studies cannot help address our research questions about the fake social media accounts that utilize GAN-generated faces. Our knowledge of these accounts is largely derived from sporadic journalistic reporting, brief communications from social media platforms, and case studies provided by hobbyists.888conspirator0.substack.com/p/a-brief-and-partial-history-of-generative We hope the present study can further our understanding of these accounts and their negative impact.
@@ 2.3 Social bots and their detection
This study also encompasses the topic of social bots, which, while related, are distinct from our primary focus. Social bots are accounts partially controlled by algorithms to emulate humans and act automatically following different instructions (Ferrara et al., 2016). To appear authentic, bots sometimes employ realistic images as profile pictures, including those stolen from other accounts999nytimes.com/interactive/2018/01/27/technology/social-media-bots.html or generated using AI models (as we show below). However, accounts using GAN-generated faces are not necessarily bots; sometimes, they are operated by real humans.
@@ 3 Data collection
We introduce multiple datasets for different purposes. The key dataset is shared with the public for further investigation and can be accessed at a public data repository.101010zenodo.org/doi/10.5281/zenodo.10436888
@@ 3.1 TwitterGAN dataset
Group name Size Date
Note
blue_check 11 2022-11-10
Subscribed to the blue-checkmark paid verification program of Twitter
bart 5 2023-02-02
Used to propagate spammy hashtags, YouTube videos, and Tumblr links
elon_watch 246 2023-02-03
Repeatedly replied to other accounts with scams involving Elon Musk
sunday 18 2023-03-05
Used to amplify the tweets of Twitter user @ilkersenock
astroturf 1,066 2023-04-08
Used to amplify content from a common set of accounts in a coordinated fashion
ind_acc 7 varies
Individual accounts that do not belong to any coordinated groups
Total 1,353
The TwitterGAN dataset comprises a total of 1,353 Twitter accounts that use GAN-generated human faces as their profiles. We collected their recent tweets using Twitter’s V2 API when it was still freely available and downloaded their profile pictures. Twitter allowed public access to user profile pictures in various resolutions. We primarily aimed for the 400-by-400-pixel versions, although some variations in resolution and aspect ratios exist.
The dataset contains multiple account groups collected between November 2022 and May 2023. The timeline of data collection and notes can be found in Table 1. The initial identification of some of these accounts was done by the Twitter user @conspirator0, while others were spotted by the authors. Following this, we conducted a thorough inspection to verify the nature of these accounts.
@@
As mentioned earlier, the GAN-generated faces are very realistic, and participants in controlled experiments can no longer distinguish them from real human faces (Nightingale and Farid, 2022). Therefore, we cannot definitely determine the nature of the accounts in TwitterGAN. However, we have identified several distinctive characteristics of GAN-generated faces that bolster confidence in their identification. First, GANs consistently show face landmarks, especially the eyes, at the same location in the generated pictures (Yang et al., 2019b). This feature becomes evident when comparing multiple GAN-generated faces side by side, as illustrated in Fig. 1. Second, many GAN-generated pictures have clear defects, which will be discussed in the next section. Third, some profile pictures contain watermarks explicitly revealing their machine-generated origin.
We also examine the profile information and activities of the accounts to triangulate their nature. For instance, there are occasional mismatches between account names and the perceived gender and race of the faces. We further rely on signals of coordinated inauthentic activities, which proved effective in similar contexts (Pacheco et al., 2021; Yang and Menczer, 2023). The TwitterGAN accounts in some groups are inter-connected through mutual following, similar profile information, and identical activity patterns. (These connections also allow us to identify more members of their groups.) Our conclusions are drawn from a holistic consideration of all these factors.
@@ 3.2 AcademicGAN dataset
Since capturing a large number of accounts using GAN-generated profiles in the wild is not trivial, we supplement the study with the AcademicGAN dataset. It contains 10,000 GAN-generated human faces obtained from the generated.photos/datasets website upon request. The website trains its own GAN models on proprietary human photos to produce more synthetic ones for sale. All pictures in AcademicGAN have the same resolution of 256 by 256 pixels and do not have the common defects observed in TwitterGAN. Nevertheless, their face landmarks are still placed at the same locations, as demonstrated in Fig. 1. This dataset is used as the ground truth in our analysis.
@@ 4.1 Profile information
Let us first focus on the profile information of the accounts in TwitterGAN. The distributions of their follower/following count, tweet count, and creation year are shown in Fig. 2. Since the astroturf group has more accounts than the other groups combined and shows distinctive features, we separate it from the rest of the accounts.
On average, accounts in astroturf have 9.95 (SD=24.03) followers, 55.68 (SD=51.75) friends, and 14.91 (SD=15.62) tweets, whereas the other accounts have 67.75 (SD=163.64) followers, 162.71 (SD=177.17) friends, and 509.23 (SD=504.02) tweets. All the differences between astroturf and the other accounts are statistically significant according to Kolmogorov–Smirnov tests ($p<0.05$). These results suggest that accounts in TwitterGAN are engaging in various activities on Twitter; however, astroturf accounts are less active than others.
Inspecting the creation years of the accounts suggests that most of them joined Twitter in the past two years, and astroturf accounts are more recent than others. This is not surprising since Twitter, both under the old management and Musk, routinely removes inauthentic accounts identified by their algorithm or reported to them. But this process is imperfect since some accounts in TwitterGAN are still alive on Twitter despite violating the policies. Specifically, all accounts in astroturf have been suspended as of December 2023, but only 14.9% accounts of the other accounts in TwitterGAN have been suspended.
@@ 4.2 Tactics and inauthentic activities
Figure 3: Illustrations of the tactics and inauthentic activities of accounts in TwitterGAN. (a) Accounts impersonating fake persons through GAN-generated profiles, human-like names, and fake descriptions. (b) Accounts replying fake screenshots with scam messages to other accounts (the keywords are redacted to avoid further spreading the scams). (c) Accounts posting spam messages repeatedly. (d) Accounts amplifying the messages from certain accounts by coordinately reposting their tweets. The accounts in red and blue represent inauthentic and normal accounts, respectively.
To uncover the tactics and inauthentic activities of accounts in TwitterGAN, we performed human annotation and computational analyses. The annotation covered all accounts from blue_check, bart, elon_watch, sunday, and ind_acc. Since the accounts in astroturf share similar patterns, we randomly selected 250 for annotation to reduce the workload. In total, we annotated 537 accounts. Two authors performed the annotation independently. The computational analyses were conducted on all accounts. Our analyses reveal several patterns, and we illustrate the four most noteworthy ones in Fig. 3. Some patterns are common across multiple groups, while others are exclusive to specific ones. We discuss the patterns next.
@@
The primary goal of using GAN-generated images in the profiles is to impersonate real humans. As illustrated in Fig. 3(a), these fake accounts often feature realistic photos, plausible names, and detailed descriptions, creating the illusion that these accounts are operated by real persons. Our analysis suggests that almost all TwitterGAN accounts employ believable names despite some being uncommon. None of these accounts disclose their use of GAN-generated profile images.
@@
Our analysis suggests that the accounts in the elon_watch group are used to promote scams. These accounts typically reply to others with screenshots depicting Elon Musk’s Twitter feed, containing deceptive messages like “Just google [keywords] and thank me later.” An illustration can be found in Fig. 3(b). The screenshots are fabricated; Elon Musk never posted these messages. Searching the keywords on Google leads to scam websites.
@@
Accounts within TwitterGAN frequently engage in spamming. We show two such examples from the bart group in Fig. 3(c), characterized by repeating messages with excessive hashtags and external links to platforms like YouTube and Tumblr.
@@
Some of the TwitterGAN accounts exhibit coordinated inauthentic activity, a tactic commonly adopted by adversarial actors during influence campaigns (Pacheco et al., 2021). A clear example is the astroturf group, where the members retweet content from a common set of accounts. The five most retweeted accounts are: @elonmusk (Elon Musk’s account), @DonaldJTrumpJr (Donald Trump Jr.’s account), @FoxNews (the official account of Fox News), @POTUS (the official account of the US President, currently owned by Joe Biden), and @Crunchyroll (the official account of crunchyroll.com, an anime distribution company). Their motives remain unclear, given the varied nature of the amplified content.
@@
Many accounts in TwitterGAN exhibit characteristics of bots. For instance, repeatedly sending scams, posting spam messages, and amplifying posts from selected accounts in a coordinated fashion are typical signs of automation. We further apply the BotometerLite model (Yang et al., 2020), a widely adopted Twitter bot detection tool, to these accounts and find that 75.2% of them are labeled likely bots (using 0.5 as threshold).
@@
Under Musk’s management, the acquisition of a blue checkmark no longer requires going through a verification process. Instead, any account willing to pay a premium fee can obtain the mark to increase its perceived legitimacy. This strategy is relatively rare among TwitterGAN accounts, possibly due to the associated cost. We only noticed the 11 accounts in blue_check adopting this strategy briefly.

### v2

@@ Abstract
Recent advancements in generative artificial intelligence (AI) have raised concerns about their potential to create convincing fake social media accounts, but empirical evidence is lacking. In this paper, we present a systematic analysis of Twitter (X) accounts using human faces generated by Generative Adversarial Networks (GANs) for their profile pictures. We present a dataset of 1,420 such accounts and show that they are used to spread scams, spam, and amplify coordinated messages, among other inauthentic activities. Leveraging a feature of GAN-generated faces—consistent eye placement—and supplementing it with human annotation, we devise an effective method for identifying GAN-generated profiles in the wild. Applying this method to a random sample of active Twitter users, we estimate a lower bound for the prevalence of profiles using GAN-generated faces between 0.021% and 0.044%—around 10K daily active accounts. These findings underscore the emerging threats posed by multimodal generative AI. We release the source code of our detection method and the data we collect to facilitate further investigation. Additionally, we provide practical heuristics to assist social media users in recognizing such accounts.
@@ 1 Introduction
Recent advances in generative artificial intelligence (AI) have resulted in remarkable competencies across various domains (Cao et al., 2023). Large language models, such as ChatGPT, [^1] https://chat.openai.com can understand human language and generate realistic text in different contexts (Jakesch et al., 2023; Ye et al., 2023). Generative vision models, such as Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) and diffusion models (Song and Ermon, 2019; Ho et al., 2020), produce photorealistic images that people can no longer distinguish from human-generated images (Lago et al., 2021; Nightingale and Farid, 2022; Bray et al., 2023). State-of-the-art multimodal models, such as Google’s Gemini models (Pichai and Hassabis, 2023), can even understand visual and audio inputs and respond with text or images. This evolving landscape of generative AI models underscores a significant leap in AI technology, opening new avenues for research and applications.
These advancements also present opportunities for misuse. For instance, adversarial actors can exploit AI models to create and disseminate disinformation or conduct influence campaigns (Yamin et al., 2021; Guembe et al., 2022; Goldstein et al., 2023; Yang and Menczer, 2024), threatening the integrity of online information exchange. In this study, we focus on the fake social media accounts that leverage generative AI for deceptive purposes (Bommasani et al., 2021). In documented cases, individuals have used AI-generated human images to fabricate fake personas online, leading to deception and targeted attacks (O’Sullivan, 2020a, b). Major social media platforms, such as Facebook (Martineau, 2019) and LinkedIn (Bond, 2022), have already encountered groups of fake accounts that utilize AI-generated faces.
@@
Although there are other generative vision models, we focus on GANs for several reasons. Firstly, the introduction of GAN-based models, particularly the StyleGAN series (Karras et al., 2019, 2020), marks a major milestone in the field of generative AI models (Cao et al., 2023) because they are capable of producing photorealistic images almost indistinguishable from real ones (Nightingale and Farid, 2022). Secondly, GANs are easily accessible. With their source code publicly available online, [^3] See https://github.com/NVlabs/stylegan and https://github.com/NVlabs/stylegan2. adversarial actors can easily produce fake human faces as needed. Additionally, for those without technical expertise, websites like https://thispersondoesnotexist.com offer an easy means to download GAN-generated images in large quantities. Last but not least, GAN-generated faces share some distinctive features, allowing us to identify them with high confidence.
In this study, we create a novel dataset of 1,420 inauthentic Twitter accounts, each utilizing a GAN-generated face as its profile picture. Upon close examination, these accounts typically feature human-like names suggesting a deliberate effort to mimic real users. These accounts engage in inauthentic activities, such as disseminating spam, executing scams, and amplifying messages in a coordinated fashion, often through automation. We further estimate the prevalence of such accounts on the platform. Leveraging the feature that GANs consistently place eyes at the same locations, together with human annotation, we develop a simple yet effective method to identify GAN-generated profiles in the wild. Applying this method to a random sample of 254,275 active Twitter users, our analysis indicates that the prevalence of accounts using GAN-generated faces is between 0.021% and 0.044%. Since our method emphasizes precision and might miss accounts leveraging more advanced AI models to generate their profiles, these estimates should be only regarded as lower bounds.
@@ 2.1 GAN-generated face detection
Although we also need to detect GAN-generated faces in this study, our case differs from previous studies significantly. The aforementioned methodology papers typically aim to distinguish generated and real human faces from datasets such as FFHQ (Karras et al., 2019) and CelebA (Liu et al., 2015). On the other hand, our objective is to identify generated faces among random Twitter profiles, which presents unique challenges. Firstly, only a subset of Twitter profiles feature human faces, and their styles may differ greatly from GAN-generated ones (see Figure 1(c) for examples). Secondly, the instances of GAN-generated faces may be relatively scarce, necessitating a detection method with high accuracy to address the class imbalance issue and provide a reliable estimate of the prevalence of fake accounts. Given these differences, existing methods are not suitable for our purpose.
@@ 2.2 GAN-generated faces on social media
Unfortunately, these studies cannot help address our research questions about the fake social media accounts that utilize GAN-generated faces. Our knowledge of these accounts is largely derived from sporadic journalistic reporting, brief communications from social media platforms, and case studies provided by hobbyists (Norteno, 2023). We hope the present study can further our understanding of these accounts and their negative impact.
@@ 2.3 Social bots and their detection
This study also encompasses the topic of social bots, which, while related, are distinct from our primary focus. Social bots are accounts partially controlled by algorithms to emulate humans and act automatically following different instructions (Ferrara et al., 2016). To appear authentic, bots sometimes employ realistic images as profile pictures, including those stolen from other accounts (Confessore et al., 2018) or generated using AI models (as we show below). However, accounts using GAN-generated faces are not necessarily bots; sometimes, they are operated by real humans.
@@ 3 Data collection
We introduce multiple datasets for different purposes. The key dataset is shared with the public for further investigation and can be accessed at a public data repository. [^4] https://zenodo.org/doi/10.5281/zenodo.10436888
@@ 3.1 TwitterGAN dataset
Group name Size Date Note
blue_check 11 2022-11-10 Subscribed to the blue-checkmark paid verification program of Twitter
bart 5 2023-02-02 Used to propagate spammy hashtags, YouTube videos, and Tumblr links
elon_watch 246 2023-02-03 Repeatedly replied to other accounts with scams involving Elon Musk
sunday 18 2023-03-05 Used to amplify the tweets of Twitter user @ilkersenock
astroturf 1,066 2023-04-08 Used to amplify content from a common set of accounts in a coordinated fashion
chatgpt 67 2024-01-31 Used ChatGPT to generate human-like content
ind_acc 7 varies Individual accounts that do not belong to any coordinated groups
Total 1,420
The TwitterGAN dataset comprises a total of 1,420 Twitter accounts that use GAN-generated human faces as their profiles. We collected their profile information and recent tweets for analysis. When Twitter’s V2 API was still available, we used it to collect the data for most of the accounts. After Twitter shut down its free API, we manually collected the data through Twitter’s web interface. We also downloaded their profile pictures. Twitter allowed public access to these pictures in various resolutions. We primarily aimed for the 400-by-400-pixel versions, although some variations in resolution and aspect ratios exist.
The dataset contains multiple account groups collected between November 2022 and January 2024. The timeline of data collection and notes can be found in Table 1. The initial identification of some of these accounts was done by the Twitter user @conspirator0, while others were spotted by the authors. Following this, we conducted a thorough inspection to verify the nature of these accounts.
@@
As mentioned earlier, the GAN-generated faces are very realistic, and layman participants in controlled experiments can no longer distinguish them from real human faces (Nightingale and Farid, 2022). However, after reading extensive literature and reviewing a large number of GAN-generated human faces, we have identified several distinctive characteristics of these faces that bolster our confidence in their identification. First, GANs consistently show face landmarks, especially the eyes, at the same location in the generated pictures (Yang et al., 2019b). This feature becomes evident when comparing multiple GAN-generated faces side by side, as illustrated in Figure 1. Second, many GAN-generated pictures have clear defects, which will be discussed in Section 4.3. Third, some profile pictures contain watermarks explicitly revealing their machine-generated origin.
We also examine the profile information and activities of the accounts to triangulate their nature. For instance, there are occasional mismatches between account names and the perceived gender and race of the faces. We further rely on signals of coordinated inauthentic activities, which proved effective in similar contexts (Pacheco et al., 2021; Yang and Menczer, 2024). The TwitterGAN accounts in some groups are inter-connected through mutual following, similar profile information, and identical activity patterns. (These connections also allow us to identify more members of their groups.) Our conclusions are drawn from a holistic consideration of all these factors.
@@ 3.2 AcademicGAN dataset
Since capturing a large number of accounts using GAN-generated profiles in the wild is not trivial, we supplement the study with the AcademicGAN dataset. It contains 10,000 GAN-generated human faces obtained from the https://generated.photos/datasets website upon request. The website trains its own GAN models on proprietary human photos to produce more synthetic ones for sale. All pictures in AcademicGAN have the same resolution of 256 by 256 pixels and do not have the common defects observed in TwitterGAN. Nevertheless, their face landmarks are still placed at the same locations, as demonstrated in Figure 1. This dataset is used as the ground truth in our analysis to compare with TwitterGAN and develop our detection method.
@@ 4.1 Profile information
Let us first focus on the profile information of the accounts in TwitterGAN. The distributions of their follower/following count, tweet count, and creation year are shown in Figure 2. Since the astroturf group has more accounts than the other groups combined and shows distinctive features, we separate it from the rest of the accounts.
On average, accounts in astroturf have 9.95 (Standard Deviation SD=24.03) followers, 55.68 (SD=51.75) friends, and 14.91 (SD=15.62) tweets, whereas the other accounts have 67.75 (SD=163.64) followers, 162.71 (SD=177.17) friends, and 509.23 (SD=504.02) tweets. All the differences between astroturf and the other accounts are statistically significant according to Kolmogorov–Smirnov tests ($p<0.05$). These results suggest that accounts in TwitterGAN are engaging in various activities on Twitter; however, astroturf accounts are less active than others.
Inspecting the creation years of the accounts suggests that most of them joined Twitter in the past two years, and astroturf accounts are more recent than others. This is not surprising since Twitter, both under the old management and Musk, routinely removes inauthentic accounts identified by their algorithm or reported to them. But this process is imperfect since some accounts in TwitterGAN are still alive on Twitter despite violating the policies. Specifically, all accounts in astroturf appeared to be suspended when we checked their status in December 2023, but only 14.9% of the other accounts in TwitterGAN had been suspended.
@@ 4.2 Tactics and inauthentic activities
Figure 3: Illustrations of the tactics and inauthentic activities of accounts in TwitterGAN. (a) Fake accounts impersonating persons that do not exist through GAN-generated profiles, human-like names, and fake descriptions. (b) Accounts replying fake screenshots with scam messages to other accounts (the keywords are redacted to avoid further spreading the scams). (c) Accounts posting spam messages repeatedly. (d) Accounts amplifying the messages from certain accounts by coordinately reposting their tweets. (e) Accounts leveraging ChatGPT to produce human-like replies. The account illustrations colored in red and blue represent inauthentic and normal ones, respectively.
To uncover the tactics and inauthentic activities of accounts in TwitterGAN, we performed human annotation and bot detection. The annotation covered all accounts from blue_check, bart, elon_watch, sunday, chatgpt, and ind_acc. Since the accounts in astroturf share similar patterns, we randomly selected 250 for annotation to reduce the workload. In total, we annotated 604 accounts, and each account was annotated by two authors independently. Our analyses reveal several patterns, and we illustrate the five most noteworthy ones in Figure 3. Some patterns are common across multiple groups, while others are exclusive to specific ones. We discuss the patterns next.
@@
The primary goal of using GAN-generated images in the profiles is to impersonate real humans. As illustrated in Figure 3(a), these fake accounts often feature realistic photos, plausible names, and detailed descriptions, creating the illusion that these accounts are operated by real persons. Our analysis suggests that almost all TwitterGAN accounts employ believable names despite some being uncommon. None of these accounts disclose their use of GAN-generated profile images.
@@
Our analysis suggests that the accounts in the elon_watch group are used to promote scams. These accounts typically reply to others with screenshots depicting Elon Musk’s Twitter feed, containing deceptive messages like “Just google [keywords] and thank me later.” An illustration can be found in Figure 3(b). The screenshots are fabricated; Elon Musk never posted these messages. Searching the keywords on Google leads to scam websites.
@@
Accounts within TwitterGAN frequently engage in spamming. We show two such examples from the bart group in Figure 3(c), characterized by repeating messages with excessive hashtags and external links to platforms like YouTube and Tumblr.
@@
Some of the TwitterGAN accounts exhibit coordinated inauthentic activity (see Figure 3(d)), a tactic commonly adopted by adversarial actors during influence campaigns (Pacheco et al., 2021). A clear example is the astroturf group, where the members retweet content from a common set of accounts. The five most retweeted accounts are: @elonmusk (Elon Musk’s account), @DonaldJTrumpJr (Donald Trump Jr.’s account), @FoxNews (the official account of Fox News), @POTUS (the official account of the US President, currently owned by Joe Biden), and @Crunchyroll (the official account of https://crunchyroll.com, an anime distribution company). Their motives remain unclear, given the varied nature of the amplified content.
ChatGPT for content generation.
Accounts in chatgpt were found to leverage ChatGPT to generate human-like content (see Figure 3(e)). They were identified by the self-revealing tweets posted by these accounts, which contained phrases such as “as an AI language model.” This method was used to uncover a larger botnet with the same tactic earlier (Yang and Menczer, 2024).
@@
Many accounts in TwitterGAN exhibit characteristics of bots. For instance, repeatedly sending scams, posting spam messages, and amplifying posts from selected accounts in a coordinated fashion are typical signs of automation. We further apply the BotometerLite model (Yang et al., 2020), a widely adopted Twitter bot detection tool, to all accounts in TwitterGAN and find that 75.2% of them are labeled likely bots (using 0.5 as threshold).
@@
Under Musk’s management, the acquisition of a blue checkmark no longer requires going through a verification process. Instead, any account willing to pay a premium fee can obtain the mark to increase its perceived legitimacy. This strategy is not very common among TwitterGAN accounts, possibly due to the associated cost. Only the 11 accounts in blue_check and 44 accounts in chatgpt adopted this strategy.